OpenStudy (anonymous):

one more : 17-(6x+3)=-6 can you explain the steps ? Thank you!

OpenStudy (anonymous):

x=10/3 do you see how i did that, using the distributive property??

OpenStudy (anonymous):

uh .. no what would be the first step?

OpenStudy (anonymous):

17-6x-3 =-6 add 6x and 6 to both sides 17+6x-6x-3+6 = -6+6+6x simplify 17-3+6 = 6x 20 = 6x or 6x =20 divide all through by 6 x =20/6 x=10/3

OpenStudy (anonymous):

is 10/3 = to 5 . becuase it says the answer should be 5 .